find the length of the hypotenuse of a right triangle with legs measuring 6 inches and 8 inches. using pythagorean theorem

Answer by stanbon(75887)

find the length of the hypotenuse of a right triangle with legs measuring 6 inches and 8 inches. using pythagorean theorem
---
h^2 = 6^2 + 8^2
h = sqrt(36+64)
h = 10 inches
======================
Cheers,
Stan H.
======================